OpenStudy (anonymous):

| x -1| -2 =| 2x -5 | solve

OpenStudy (anonymous):

please help me

OpenStudy (nincompoop):

there are two things to remember when dealing with absolute values that the x can be less than zero or greater than zero meaning you would have either a positive value for it or a negative value for it.

OpenStudy (anonymous):

make 4 cases of x>=1 x<1, x>=5/2, x<5/2

OpenStudy (anonymous):

and I solve for x ??

OpenStudy (nincompoop):

naturally, that is what you do :D

OpenStudy (anonymous):

can you just give me the first equation plz

OpenStudy (anonymous):

Guys !1

hartnn (hartnn):

ok, we use the simple fact that when |a|=b , either a=b or a=-b we need to use that twice here, because we have |....| twice. so see this example : |a|+b=|c| gives 2 equations, (|a|+b=c) or (|a|+b=-c) which again gets split to 4 equations, (a+b=c or -a+b=c) or (a+b=-c or -a+b=-c) read this carefully and tell me whether u get it...

hartnn (hartnn):

ask if u have any doubt in any step.

OpenStudy (anonymous):

ok I'll do my best

OpenStudy (skullpatrol):

nice work @hartnn

hartnn (hartnn):

hey thanks :)

hartnn (hartnn):

so @hataki lets first split this into 2 equations, | x -1| -2 =| 2x -5 | which 2 equations ?

OpenStudy (anonymous):

| x -1| -2 =| 2x -5 | or | x -1| -2 = -| 2x -5 | right ?

hartnn (hartnn):

when you have splitted, no need to use |...| again! so its | x -1| -2 = 2x -5 or | x -1| -2 = -( 2x -5 ) ok ?

OpenStudy (anonymous):

ok get it and now we split them into 4 equations

hartnn (hartnn):

yes, the idea is same and simple., |a| = b gives a=b or a=-b so, lets first only take 1st equation and split it into 2 | x -1| -2 = 2x -5 2 equations ? (if you are confident, write all 4 equations :))

OpenStudy (anonymous):

1) x-1-2=2x-5 2) -x+1-2=2x-5 3)x-1-2=-(2x-5) 4)-x+1-2=-(2x-5) right ??

hartnn (hartnn):

oh, wow , absolutely correct! :) you learn fast.

hartnn (hartnn):

just solve them individually and get 4 values of x first....they are simple algebraic equations with no twists now...

OpenStudy (anonymous):

THANK YOU SO SO SO MUCH :D if I messed up can I get back to you ??

OpenStudy (anonymous):

ANd thanks to every one :D

hartnn (hartnn):

sure, please tell us what you get finally, we will verify it! also, its not always necessary that all 4 values are solution, we need to cross verify, so you better tell us what 4 values u get, then we will proceed on verification

OpenStudy (anonymous):

ok I'm work on them NOW :D

OpenStudy (anonymous):

1) x=2 2) x=4/3 3) x=8/3 4) x=6

hartnn (hartnn):

all correct! :)

hartnn (hartnn):

now we put these one by on in original equation! put x=2 in | x -1| -2 =| 2x -5 | |1| -2 = |-1| so, 1 - 2 = 1 which is false! so, x= 2 is not the solution, got this ?

OpenStudy (anonymous):

yeah I'm working on them :D

OpenStudy (anonymous):

sir non of them worked they all not equal

hartnn (hartnn):

thats absolutely correct! there is actually no solution to your original equation :)

OpenStudy (anonymous):

thank you so much sir I do understand it now :D :D

hartnn (hartnn):

you are welcome ^_^